OpenStudy (taramgrant0543664):

Noticed a few questions lately about aromaticity and I needed a small reminder so I put a little tutorial like thing! Let me know if there's anything I missed or if I should clarify anything! Thanks

OpenStudy (taramgrant0543664):

Aromaticity is a property of conjugated cycloalkenes in which the stabilization of the molecule is strong due to the ability of the electrons in the pi orbitals to delocalize and act as a framework to create a planar molecule. The three general requirements for a compound to be aromatic are: The compound must be cyclic Each element within the ring must have a p-orbital that is perpendicular to the ring, hence the molecule is planar. The compound must follow Hückel's Rule (the ring has to contain 4n+2 p-orbital electrons). The simplest compound that is aromatic is benzene. Benzene (C6H6) is a cyclic compound, it is planar and follows Huckel's Rule. It follows Huckel's Rule because 4n+2=6 (as benzene has 6 pi electrons) So n=1, which is a positive integer, and therefore follows the rule. Note: Pi electrons are sp2 hybridized

OpenStudy (taramgrant0543664):

Example 1:

OpenStudy (taramgrant0543664):

Example 1: This example is not aromatic because it doesn't follow Huckel's Rule as the top carbon is sp3 hybridized.

OpenStudy (taramgrant0543664):

Example 2:

OpenStudy (taramgrant0543664):

Example 2: This is aromatic as it follows the aromatic requirements and looking at Huckel's Rule n=1

OpenStudy (photon336):

Once again Tara job well done

OpenStudy (photon336):

@taramgrant0543664 I sometimes hear of the term Anti aromatic, is that the same as non-aromatic?

OpenStudy (taramgrant0543664):

Awe that's a great question! Never even thought about adding that in. So aromatic compounds An aromatic molecule is: Flat, Cyclic, Completely conjugated and the number of π electrons must obey Huckel’s rule of 4n+2. An anti-aromatic molecule is much like the aromatic compound in that it is also Flat, Cyclic, and Completely conjugated but the difference is that the number of π electrons follows the formula of 4n instead of the 4n+2. A non-aromatic molecule is: NOT flat, NOT cyclic, and NOT completely conjugated. Some examples include the typical alkane and/or alkene such as butane or hexane.

OpenStudy (taramgrant0543664):

Example 3: is anti aromatic. It is cyclic, it is fully conjugated but it doesn't follow Huckel's Rule of 4n+2 it follows the 4n rule 4n=4 (4 pi electrons) so n=1

OpenStudy (photon336):

Sometimes you have an aromatic system, I.e a heterocyclic compound, with a (oxygen or nitrogen atom in the center) sometimes those lone pair electrons participate in the conjugated system; if something is like your example #2 but does that nitrogen have any pi electrons? because that's sp^3 hybridized so those electrons would have to be in a p orbital I guess

OpenStudy (photon336):

Thanks for that clarification